# Retrieval Augmented Instruction Tuning for Open NER with Large Language Models

### v1

@@ 3.1 Experimental Settings
Backbones: We adopt LLaMA-3-8B (Meta, 2024) and Qwen-1.5-7B (Team, 2024) as the backbone models for English and Chinese scenarios respectively. Training data: For English, we use the training data Pile-NER released by Zhou et al. (2024). For Chinese, we use the training data Sky-NER constructed in this paper as described in Section 3.2. We use LoRA (Hu et al., 2021) to train models. Retrieval: We adopt GTE-large [^2] https://huggingface.co/thenlper/gte-large (Li et al., 2023) to generate text embeddings and set $k=2$ in main experiments. Evaluation: We mainly focus on the zero-shot evaluation. For English, we adopt benchmarks CrossNER, MIT-Movie and MIT-restaurant following Zhou et al. (2024). For Chinese, we collect eight benchmarks across diverse domains, of which details are in Appendix D. We report micro-F1 value.
@@ 3.2 Chinese IT Data Construction
Following the data construction recipe of UniNER (Zhou et al., 2024), we construct an IT dataset for Chinese open NER. We sample input passages from the large-scale Sky corpus (Wei et al., 2023) across various domains, then use ChatGPT (gpt-3.5-turbo) to generate entity mentions and types based on the sampled passages. More details of data construction procedures are in Appendix B. We name this dataset as Sky-NER, which consists of 50K NER examples, and the type statistics are in Table 3.
@@ 3.5 Analysis
Figure 3: Impacts of training using various retrieval strategies. The average F1 value of the evaluated benchmarks are reported. NN shows the best performances, suggesting the need of training with retrieved context.
@@
Figure 5: Impacts of inference with in-domain examples. The average F1 value of the evaluated benchmarks are reported. $N$-exmp. means the example pool of size $N$. Sufficient in-domain examples are helpful for RAG.
@@ 5 Limitations
Appendix A Related Work
A.1 IE with LLMs
The main techniques studied in the area of IE with LLMs fall under advanced prompt designing(Guo et al., 2023; Xie et al., 2023; Wang et al., 2023a), instruction tuning (IT) (Sainz et al., 2024; Zhou et al., 2024; Li et al., 2024) and data augmentation (Josifoski et al., 2023; Zhang et al., 2023; Ma et al., 2023). Many of the prompt designing methods apply RAG to an off-the-shelf LLM to assist inference (Guo et al., 2023; Wan et al., 2023; Xie et al., 2024), which retrieves similar examples to provide more useful information the LLM. Works of IT incorporate the information for IE into the LLMs through task-specific fine-tuning (Sainz et al., 2024; Zhou et al., 2024). Different from previous works, we explore retrieval augmented IT (RA-IT) for IE, with a focus on the open NER task.
Following UniNER (Zhou et al., 2024), we conduct investigations under the targeted distillation setting, since UniNER successfully distills the strong capability of ChatGPT in open NER into a smaller model without any human-annotated data. Other works of IT for IE, Sainz et al. (2024); Li et al. (2024) adopt the code-style instruction to fine-tune LLMs in effectively generating IE outputs through code generation. They are orthogonal to this work since the strategy of RA-IT can be integrated to various styles of instructions.
A.2 Retrieval aware Fine-Tuning
Appendix B Chinese Data Construction
@@
Appendix C Diverse Teachers for Data Construction
@@
Appendix D More Details of Experimental Settings
D.1 Preliminary Study on Data Efficiency
@@ D.4 Benchmarks
Appendix E Full results on data size study
@@
Appendix F Extended Analysis
F.1 Varying Numbers of Examples
@@ F.2 Ablation for Mixed NN
Appendix G Examples of Instruction Data

### v2

@@ 3.1 Experimental Settings
Backbones: We adopt LLaMA-3-8B (Meta, 2024) and Qwen-1.5-7B (Team, 2024) as the backbone models for English and Chinese scenarios respectively. Training: For English, we use the training data Pile-NER released by Zhou et al. (2024). For Chinese, we use the training data Sky-NER constructed in this paper as described in Section 3.2. We use LoRA (Hu et al., 2021) to train models. Our training infrastructure was 1 NVIDIA A100 80GB. Retrieval: We adopt GTE-large [^2] https://huggingface.co/thenlper/gte-large (Li et al., 2023) to generate text embeddings and set $k=2$ in main experiments. Evaluation: We mainly focus on the zero-shot evaluation. For English, we adopt benchmarks CrossNER, MIT-Movie and MIT-restaurant following Zhou et al. (2024). For Chinese, we collect eight benchmarks across diverse domains, of which details are in Appendix C. We report micro-F1 value.
@@ 3.2 Chinese IT Data Construction
Following the data construction recipe of UniNER (Zhou et al., 2024), we construct an IT dataset for Chinese open NER. We sample input passages from the large-scale Sky corpus (Wei et al., 2023) across various domains, then use ChatGPT (gpt-3.5-turbo) to generate entity mentions and types based on the sampled passages. More details of data construction procedures are in Appendix A. We name this dataset as Sky-NER, which consists of 50K NER examples, and the type statistics are in Table 3.
@@ 3.5 Analysis
Figure 3: Impacts of training using various retrieval strategies in RA-IT. The average F1 value of the evaluated benchmarks is reported. NN exhibits the best performances, suggesting the need of training with retrieved context.
Figure 5: Impacts of inference with in-domain examples using NN retrieval. The average F1 value of the evaluated benchmarks are reported. $N$-exmp. means the example pool of size $N$. The results indicate that sufficient in-domain examples are helpful for inference with RAG.
@@
4 Related Work
4.1 IE with LLMs
The main techniques studied in the area of IE with LLMs fall under advanced prompt designing(Guo et al., 2023; Xie et al., 2023; Wang et al., 2023a), instruction tuning (IT) (Sainz et al., 2024; Zhou et al., 2024; Li et al., 2024) and data augmentation (Josifoski et al., 2023; Zhang et al., 2023; Ma et al., 2023). Many of the prompt designing methods apply RAG to an off-the-shelf LLM to assist inference (Guo et al., 2023; Wan et al., 2023; Xie et al., 2024), which retrieves similar examples to provide more useful information for the LLM. Works of IT incorporate the information for IE into the LLMs through task-specific fine-tuning (Sainz et al., 2024; Zhou et al., 2024). Different from previous works, we explore retrieval augmented IT (RA-IT) for IE, with a focus on the open NER task.
Following UniNER (Zhou et al., 2024), we conduct investigations under the targeted distillation setting, since UniNER successfully distills the strong capability of ChatGPT in open NER into a smaller model without any human-annotated data. Other works of IT for IE, Sainz et al. (2024); Li et al. (2024) adopt the code-style instruction to fine-tune LLMs in effectively generating IE outputs through code generation. They are orthogonal to this work since the strategy of RA-IT can be integrated into various styles of instructions. Moreover, Zaratiana et al. (2023) integrated the strong capability of ChatGPT in open NER into smaller-scale bidirectional LMs (BiLMs) such as BERT (Devlin et al., 2019). How to integrate retrieval augmentation into the BiLMs frameworks is also worth exploring in future work.
4.2 Retrieval aware Fine-Tuning
@@ 5 Conclusion
Acknowledgements
This research is supported by Zhejiang Provincial Natural Science Foundation of China (LDT23F02023F02) and the National Natural Science Foundation of China (72350710798).
Appendix A Chinese Data Construction
@@
Appendix B Diverse Teachers for Data Construction
@@
Appendix C More Details of Experimental Settings
C.1 Preliminary Study on Data Efficiency
@@ C.4 Benchmarks
Appendix D Full results on data size study
@@
Appendix E Extended Analysis
E.1 Varying Numbers of Examples
@@ E.2 Ablation for Mixed NN
Appendix F Examples of Instruction Data
@@
Case 1
Input text What is the theme song to stand by me?
Ground Truth [{’stand by me’: ’title’}]
Answer from Vanilla-IT [{’stand by me’: ’song’}, {’theme song’: ’song’}]
Answer from RA-IT [{’stand by me’: ’title’}, {’theme song’: ’song’}]
Case 2
Input text How many times has matt damon been jason bourne?
Ground Truth [{jason bourne: character}, {matt damon: actor}]
Answer from Vanilla-IT [{paul greengrass: director},{jason bourne: title}, {matt damon: actor}]
Answer from RA-IT [{jason bourne: character}, {matt damon: actor}]
Table 9: Cases that RA-IT benefits the long-tail entity types. We exclude the top 30% of frequent types and regard the remaining types as long-tail types. The entities in bold are long-tail types that are misclassified by vanilla-IT and corrected by RA-IT. These two cases are also commonsense-related.
Case 1
Input text Viral TK phosphorylates aciclovir into its monophosphate form , which is subsequently phosphorylated to active aciclovir triphoshate by cellular kinase s, thus selectively inhibiting viral DNA polymerase.
Ground Truth [{Viral TK: enzyme}, {cellular kinase: enzyme}, {DNA polymerase: enzyme}, {aciclovir triphoshate: chemical compound}]
Answer from Vanilla-IT [{Viral TK: scientist}, {cellular kinase s: scientist}, …]
Answer from RA-IT [{Viral TK: scientist}, {cellular kinase s: scientist}, …]
Case 2
Input text NIST also differs from Bilingual evaluation understudy in its calculation of the brevity penalty insofar as small variations in translation length do not impact the overall score as much.
Ground Truth [{NIST: metrics}, {bilingual evaluation understudy: metrics}]
Answer from Vanilla-IT [{NIST: organization}, {bilingual evaluation understudy: organization}, …]
Answer from RA-IT [{NIST: organization}, {bilingual evaluation understudy: organization}, …]
Table 10: Cases where RA-IT fails to improve. The entities in bold are vanilla-IT wrongly recognized, and RA-IT failed to improve. These professional entities in biomedical or AI domains require domain knowledge to be recognized.
Appendix G Case Study
We conduct case study to explore the advantages and disadvantages of the proposed RA-IT method. Table 9 are two cases that demonstrate that RA-IT benefits the long-tail entity types. We exclude the top 30% of frequent types and regard the remaining types as long-tail types. The entities in bold are long-tail types that are misclassified by vanilla-IT and corrected by RA-IT. These two cases are also commonsense-related.
Table 10 are some bad cases where RA-IT fails to improve. The entities in bold are vanilla-IT wrongly recognized, and RA-IT failed to improve. These professional entities in biomedical or AI domains require domain knowledge to be recognized. This shows that RA-IT benefits commonsense-related cases more than knowledge-seeking cases.